OpenStudy (anonymous):

Solve for x. -ax + 3 b > 5

OpenStudy (anonymous):

-ax + 3b > 5 -3b -3b -ax > 5 - 3b divided everything by -a x < (-5/a) + (3b/a)

OpenStudy (anonymous):

If you want it in as one fraction: x < (3b-5)/a

OpenStudy (anonymous):

Now, I can't I have 3b in the front of 5?

OpenStudy (anonymous):

Plus, there isn't an answer for that one. I got: x < 3b + 5 over a

OpenStudy (anonymous):

(3b/a)-(5/a) = (-5/a)+(3b/a) = (-5+3b)/a = (3b-5)/a

OpenStudy (anonymous):

So...the answer is my answer?

OpenStudy (anonymous):

3b-5 is all divided by a, not just 5

OpenStudy (anonymous):

I know, but there isn't a choice for that answer. :/

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

-ax+3b>5 -3b -3b -------------- -ax>5-3b --- ----- -a -a x<5-3b ------ -a

OpenStudy (nikato):

this is what i got

OpenStudy (anonymous):

Same thing as me.

OpenStudy (anonymous):

Ok, I'm trusting your answers you guys. Thank you.

OpenStudy (anonymous):

(-3b/-a) -> (3b/a)

OpenStudy (nikato):

yea. no problem

OpenStudy (anonymous):

Wait

OpenStudy (nikato):

yes?

OpenStudy (anonymous):

That's not a solution. Here are the solutions: x > the quantity 3 times b minus 5 all over negative a x > the quantity 5 minus 3 times b all over negative a x < the quantity 3 times b plus 5 all over negative a x < the quantity negative 3 times b plus 5 all over negative a

OpenStudy (anonymous):

I mean C.

OpenStudy (nikato):

i think its D

OpenStudy (anonymous):

why

OpenStudy (anonymous):

D is what we both got as the answer.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

all over -a is (3b-5)/a

OpenStudy (anonymous):

Ok, I'm going w/ your answer D.; thank you both.

OpenStudy (anonymous):

*all over -a is (3b-5)/-a also: http://www.wolframalpha.com/ This website solves any equation.

OpenStudy (anonymous):

So, you're telling me that the answer is D. and I love that website.